The equation for the marginal productivity of capital is given​ by

MPK-1,000 10 The price of a unit of capital is 2,000 The rate of depreciation is: 5% per year. The real interest rates: 9% per year If the existing level of capital Kt is equal to 60 units, what is the level of gross investment? It = units (enter your answer rounded to one decimal place)
Answer #1

User cost of capital = ( r+d ) * P_k

=( .09+.05)*2000

= 280

r : interest rate, d: depreciation rate, P_k : price of capital

At desired level of capital stock :

MPK = user cost of capital

1000-10*K = 280

Thus K* = 72   

Thus gross level of investment = K* - current capital stock

= 72-60 = 12

Thus I_t = 12 units answer